The equilibrium constant for the chemical equation N2 g) +3H 2(g) 근 2NH3(g) is Kp = 71.7 at 177 °C. Calculate the value of the Ke for the reaction at 177 °C. Number K,= Tools x 102
Answer #1

\Delta n = 2 - (1+3) = - 2

K_p=K_C \times (RT)^{ \Delta n }

K_C = \dfrac { K_p} { \times (RT)^{ \Delta n }}

K_C = \dfrac { 71.7} { \times (0.08206 \: L atm /mol/K \times (177+273.15))^{ -2 }}

K_C =97836

To three significant figures, the answer is 97800